OpenStudy (anonymous):

Find the point(s) where the line tangent to the graph of f(x)= 2x^3+6x^2-18x-2 is horizontal

OpenStudy (anonymous):

I found the first deriv

OpenStudy (anonymous):

the slope, but stuck on trying to find points

OpenStudy (tad1):

If the tangent to the graph is horizontal, the first derivative, which is the slope , = 0. So the equation will be y = a constant. So set the first derivative = 0 6x^2 + 12x - 18 = 0 factor out the 6 to get 6(x^2 + 2x - 3) = 0 Factor (x-3)(x+1) = so X = 3 or X = -1 Substitute into f(x) to find y.

OpenStudy (anonymous):

oh of course! horizontal line has not slope! Thank you!

OpenStudy (tad1):

It does have a slope; the slope is zero. a vertical line has no slope.

OpenStudy (anonymous):

plugging in 3 though gives me the wrong y coordinate - i think it might be -3

OpenStudy (anonymous):

unless, i did something wrong

OpenStudy (anonymous):

nope should be x+3 as a factor therefore x=-3

OpenStudy (tad1):

You're right. Sorry for the error. Tad

OpenStudy (anonymous):

no worries, thanks again for the help

OpenStudy (tad1):

Thanks for the medal.

OpenStudy (radar):

I got two points for y(-3)=52 for a point (-3, 52) and Y(1)=-12 for a point (1, -12) Now which point is correct or is there two points?

OpenStudy (radar):

I noticed that the problem asked for point(s), leaving that as a possibility.

OpenStudy (tad1):

There are two points. where the slope of the tangent line = 0. the original function f(x) is cubic. It would have three roots, though two of them may be complex.

OpenStudy (tad1):

The degree of the equation tells you how many roots it has, though sometimes the roots are repeated as in a perfect square quadratic.

OpenStudy (radar):

OK. so I guess the horizontal lines would be: y=-12 and y=52 is that what you came up with?